OpenStudy (anonymous):

Solve square root of x-9 minus square root of x = -1

OpenStudy (anonymous):

\[\sqrt{x-9}-\sqrt{x}=-1\] \[\sqrt{x-9}\ge0,x \ge9\] put \[\sqrt{x-9}=a,\sqrt{x}=b,a-b=-1\] \[a^2-b^2=x-9-x=-9\] \[\frac{ a^2-b^2 }{ a-b }=\frac{ -9 }{ -1 }=9,a+b=9\] adding a-b+a+b=-1+9=8 2a=8 a=4 \[\sqrt{x-9}=4,squaring,\] x-9=16 x=16+9=25

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw